OpenStudy (anonymous):

I really need help! Can someone explain this to me? I give medals and fan!!! Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (anonymous):

Here are the answer choices. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (juana02):

Y=-6x-15 top one is the answe

OpenStudy (anonymous):

Thanks! Can you explain that?

OpenStudy (juana02):

Sure

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Can you explain like right now? lol

OpenStudy (juana02):

I'm trying to attach the picture but I cant

OpenStudy (anonymous):

Oh. That's okay, thanks anyway!!

OpenStudy (juana02):

Basically the equation form of points -2 and -3 and slope -6 is the top answer

OpenStudy (juana02):

I have to type it I can't attach the picture

OpenStudy (juana02):

Do u understand now?

OpenStudy (anonymous):

Yeah, thanks!

OpenStudy (juana02):

When will u reward me with the medal? @dan815

OpenStudy (juana02):

@dancingdiamond

OpenStudy (anonymous):

I already did, and I fanned you

OpenStudy (juana02):

Thanks any other questions just tag me

OpenStudy (anonymous):

Thanks!!